AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 24 of 33  ELOG logo
ID Date Author Subject Text Attachments
  204   Thu Feb 17 13:12:19 2022 RSS, AGAlpha background run4314:00

 

Check Leakage current: (see attachment
 11x 
  203   Thu Feb 17 08:31:07 2022 RSS, AGAlpha background run4309:15

Check Leakage current: (see attachment
2)
 11x 
  202   Wed Feb 16 21:41:11 2022 JMRun 43 Alpha Background22:41

Check Leakage current: (see attachment
1) = 10.01 uA and falling
 8x 
  201   Wed Feb 16 21:27:31 2022 JMCryring signals, Pulser tests, Waveforms On 15/2/22 it was noted that the leakage
current increases when the motors are moved
IN. This is contary to tests conducted on
 11x 
  200   Wed Feb 16 15:27:53 2022 EM, JMRun 41 Pulser tests16:30

NOTE: Pulser is ON!

Check Leakage current: (see attachment
 11x 
  199   Wed Feb 16 12:43:05 2022 TP, RSSAlpha background run40 13:30

Check Leakage current: (see attachment
1)
 11x 
  198   Wed Feb 16 08:27:27 2022 TP, RSSAlpha background run4009:15

Check Leakage current: (see attachment
1)
 11x 
  197   Tue Feb 15 21:28:14 2022 JMNoisy ChannelsAll fees and asics checked to find noisy
channels. Channels are listed along with
the noise peak height for the low energy
 2022-02-15_22-17-51-aida1-3*L.png2022-02-15_22-23-12-aida1-4-stat-log.png 
  196   Tue Feb 15 14:59:37 2022 JM, CB, RJCPulser walkthrough16:00

Check Fee temperatures: all green
(see attachment 1)
 11x 
  195   Tue Feb 15 10:46:37 2022 RSS, LV, CBAlpha background run3711:30

Check Fee temperatures: all green
(see attachment 1)
 10x 
  194   Mon Feb 14 09:56:39 2022 CBShifters instructions 2022EMERGENCY SHUTDOWNS

1. CARME motors: Q key
on keyboard while on workspace 1 of appc218
 10x 
  193   Sun Feb 13 21:28:35 2022 JM, CBAlpha background Run 32, 35Alpha background gathered from 17:35 12/2/22
to 13:44 13/2/22 is run32. Run time ~20 hours.
Some alphas visible in aida1 1*L histograms
 6x 
  192   Sat Feb 12 19:48:38 2022 JM, CBGrounding and movement tests> Initial grounding see final grounding
from elog 187. ASIC settings for all - LEC
slow comparator 0x64. Note on 11/02/22 1.8*L
 11x 
  191   Thu Feb 10 09:12:41 2022 OH, CB, USHourly motor movementTo test effect of the motors on the leakage
current, they were cycled once per hour overnight
at approximately 50 minutes past the hour.
 220209_1014_leakage.png220210_1443_BiasFeeOn.png 
  190   Wed Feb 9 08:01:41 2022 OH, CB, G-HCResults from no beam overnight08:52 Overnight the DAQ was left running to
check if there was any individual radioactivity
in the Si.
 220209_0852_leakage.png220209_0900_leakage_shake.png 
  189   Tue Feb 8 08:13:53 2022 OH, CB, DFHopefully effect of beam tuning on detectors09:00 Detector still biased to 60V and leakage
current has been mostly stable overnight.
      Only minor fluctuations observed -
 9x 
  188   Mon Feb 7 08:09:50 2022 OH, CB and PHCARME Server setup and initial tests09:09 Arrived at GSI for comissioning beamtime.
      CARME currently in a stopped state,
with all FEEs powered off.
  
  187   Fri Feb 4 12:49:54 2022 TD, JMGrounding Tests and current status
JM plans to install new crimped, 3mm dia
copper drain wires (a la AIDA@DESPEC) this
 10x 
  186   Fri Feb 4 09:53:48 2022 JM, TD, OH (remotely)Alpha spectra overnightDAQ was left running last night from 19:30.
Detector bias 80V with a leakage current
~3.4 uA. Pulser OFF.
 8x 
  185   Thu Feb 3 22:05:33 2022 JM, TD, OH(remotely)Bias update and grounding9:30 Bias remained stable overnight - 60V
2.95 uA.

10:00 No grounding attached initailly.
  
ELOG V3.1.4-unknown